Shamrock Office Supplies is interested in estimating the cost involved in hiring new employees. The following information is available regarding the costs of operating the Human Resource department at Shamrock Office Supplies in May, when there were 64 new hires:

Human Resources Department

May

Staff wages

$31,360

Manager salary

7,240

Office supplies

320

Depreciation of office equipment

440

Share of building cost (based on square feet
  occupied by Human Resources)

1,540

Total

$40,900

Use account analysis to determine fixed cost per month and variable cost per new hire. Assume that only staff wages and office supplies are variable costs.

Fixed Cost

$

per month

Variable Cost

$

per new hire

The company is planning to hire 76 employees in June. Estimate the total cost of Human Resources for June.

Total cost of Human Resource $

What is the expected incremental cost associated with hiring 12 more employees than were hired in May?

Expected incremental cost $
Answer #1

A] FIXED COST PER MONTH = DEPRECIATION + BUILDING COST

= 440+1,540

=$1,980

VARIABLE COST PER NEW HIRE = (STAAF SALARIES+MANAGER SALARY+OFFICE SUPPLIES)/64

=(31,360+7,240+320)/64

=$608 PER NEW HIRE

B]76 EMPLOYEES IN JUNE

TOTAL COST OF HUMAN RESOURCES FOR JUNE = 1,980 + 608*76

=$48,188

C] The expected incremental cost associated with hiring 12 more employees than were hired in May

=[(64+12)*608 + 1,980]

=$48,188

I hope it is useful to u if u have any doubt plz comment and plz give me up thumb